OpenStudy (xxorejina):

Find the slope of a line parallel to 2y=6x+8. 3 -3 6 -6

Nnesha (nnesha):

what's the slope of given line ??

OpenStudy (shadowlegendx):

I know this one :D

Nnesha (nnesha):

good you can help him/her if you want

OpenStudy (alexandervonhumboldt2):

slopes of parallel lines are same :)

OpenStudy (anonymous):

2y=6x+8. (divide both sides by 2) y = 3x + 4 (according to y=mx+b, m=slope) slope=3 since parallel lines have same slopes the slope of a line parallel to this line would also be 3